OpenStudy (anonymous):

The ages of 5 children are 2, 4, 6, 8 and 10 years. Is the mean or median the best measure of center to summarize the ages of the children? A. mean B. median C. both

OpenStudy (anonymous):

@futureescape

OpenStudy (futureescape):

A

OpenStudy (anonymous):

well the mean means the average but the median is the middle number so I would think the best answer is C

OpenStudy (anonymous):

ahhh two different answers!

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

im not typing why does it always say im typing....

OpenStudy (anonymous):

Because you were typing that you weren't typing.

OpenStudy (anonymous):

lol

OpenStudy (futureescape):

@tayromo sounds right. C was my second choice. I think i'm changing my answer.

OpenStudy (anonymous):

there isnt a choice of D lol

OpenStudy (anonymous):

ok ill go with C I wanted to say C to but I just wasnt sure

OpenStudy (anonymous):

It's C because both the median AND mean are 6.

OpenStudy (anonymous):

\[\frac{2+4+6+8+10}{5} = \frac{30}{5} = 6 = mean = median = 6\]